### v2

# On nodal deformations of singular surfaces in $\mathbb{P}^{3}$

Abstract.
In this paper we study nodal deformations of singular surfaces $S\subset\mathbb{P}^{3}$. In particular we consider the case in which $S$ has an isolated singularity of multiplicity $m$ and the case in which $S$ has only ordinary singularities along a line.
Key words and phrases:
Severi varieties, surfaces, deformations, singularities, nodes, pinch points, nodal deformations of pinch points
1991 Mathematics Subject Classification:
14B05, 14B07, 14J10, 14J17
1. Introduction
Let $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}\subset|\mathcal{O}_{\mathbb{P}^{3}}(n)|$ be the Severi variety of surfaces $S$ of degree $n$ with $\delta\geqslant 1$ ordinary double points (called nodes) and no other singularity. Then $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ is a locally closed subvariety of $|\mathcal{O}_{\mathbb{P}^{3}}(n)|\cong\mathbb{P}^{{{n+3}\choose 3}-1}$. Of course $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{1}$ is irreducible of codimension 1 in $|\mathcal{O}_{\mathbb{P}^{3}}(n)|$ and its closure $\mathbb{S}_{n}\subset|\mathcal{O}_{\mathbb{P}^{3}}(n)|$ is the set of all singular surfaces of degree $n$. If $S\in\mathbb{S}_{n}$ is a singular surface, it makes sense to consider the maximum $\delta_{S}$ of the integers $\delta$ such that the point [S] corresponding to $S$ belongs to the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$.
In this article we will consider the following:
Problem 1.1.
Let $S\in\mathbb{S}_{n}$ be a degree $n$ singular surface in $\mathbb{P}^{3}$. In the closure of which Severi varieties $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ the point [S] corresponding to $S$ is contained? In particular, compute $\delta_{S}$.
Though this question is rather basic, it does not seem to have received much attention so far.
One can ask a similar question for plane curves, and it seems to have been considered for the first time by G. Albanese in [1], where he claimed that an irreducible plane curve of degree $n$ and genus $g$ sits in the closure of the Severi variety $V^{\mathbb{P}^{2},|\mathcal{O}_{\mathbb{P}^{2}}(n)|}_{\delta}$, with
$\delta={{n-1}\choose 2}-g$
and $\delta$ is, of course, the maximum integer for which this happens. However Albanese’s arguments do not meet the present standard of rigor, and this theorem has been proved in more recent times by Arbarello–Cornalba [2] and Zariski [12].
In the case of surfaces in $\mathbb{P}^{3}$ the problem is much more complicated and we have no idea of what could be a general answer to it. In the present paper, which should be regarded as having an experimental flavor, we focus on two study cases, namely: (a) the case in which $S$ has a single ordinary or quasi–ordinary singularity of multiplicity $m$ (for the definitions see the beginning of Section 3) and, (b) the case in which $S$ has ordinary singularities along a line, i.e., at the general point of the line $S$ has normal crossings, and at finitely many points of the line it has pinch–points (i.e., the tangent cone consists there of a plane with multiplicity 2).
Our methods are based on deformation theory of regular components of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ whose closure may contain the point [S] corresponding to $S$.
Recall that an irreducible component $V$ of a Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ is said to be regular if for the general point $S]\in V$ one has $h^{1}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n)\otimes\mathcal{I}_{N})=0$, where $N$ is the reduced scheme of lenght $\delta$ of the singular points of $S$. This means that the points in $N$ give $\delta$ independent conditions to the surfaces in $|\mathcal{O}_{\mathbb{P}^{3}}(n)|$. If $V$ is regular, then $V$ is reduced and
$\dim(V)=h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n))-1-\delta={{n+3}\choose 3}-1-\delta.$
Note that if [S] is in the closure of an irreducible regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, then it sits also in the closure of an irreducible regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta^{\prime}}$ with $\delta^{\prime}\leqslant\delta$, because the nodes in a regular component of the Severi variety can be independently smoothed.
Case (a) is considered in Section 3. In Corollary 3.5 we give an answer to Problem 1.1 for $3\leqslant m\leqslant 5$. This is the best possible answer for $m=3$. For $4\leqslant m\leqslant 5$ this is the best possible answer provided one looks only at deformations in regular components of the Severi variety. In Corollary 3.5 we also provide a partial answer for any $m$. In particular we prove that if $S\subset\mathbb{P}^{3}$ is surface of degree $n$ with a unique general ordinary singularity of multiplicity $m$, then the point corresponding to $S$ is in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{{m-1}\choose 2}$. An ordinary singularity of multiplicity $m$ is said to be general if its tangent cone is the cone over a general plane curve $C$ of degree $m$. In any case, we show that the answer to Problem 1.1 in case (a) definitely depends on the fact that the singularity is ordinary or quasi–ordinary.
In order to attack case (a) we first need to consider in Section 2 a preliminary, equally interesting problem, i.e., Problem 2.1: given an irreducible component $V$ of a Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, and given $S]\in V$ a general point and $C$ a general plane curve section of $S$, when is it the case that $C$ is a general plane curve of degree $n$? Again we are unable to solve this problem in general, but we prove some partial results, of independent interest (see, e.g., Proposition 2.4 and its consequences) that are useful to prove the aforementioned results in case (a).
Case (b) is treated in Section 5. Our main result here is Theorem 5.7, that says that if $S$ is a general surface of degree $n\geqslant 3$ that has only ordinary singularities along a line, then $S$ belongs to the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{3n-4-\epsilon}$, with $0\leqslant\epsilon\leqslant 1$ having the same parity of $n$. In particular we prove that we can choose $\ell$, with $3n-4=2\ell+\epsilon$, pinch points of $S$ on the line so that $S$ deforms to a $2\ell$–nodal surface in a regular component of the corresponding Severi variety, so that each chosen pinch point deforms to two nodes. The integer $3n-4-\epsilon$ is the maximum for which this happens in the $n$ even case, whereas in the $n$ odd case the maximum could be 3n43n-4.
To prove this result we have to rely on the existence of suitable rational surfaces of degree $n$ with a line $r$ of points of multiplicity n2n-2 and with $3n-4-\epsilon$ nodes off $r$, that impose independent conditions to surfaces of degree $n$ in $\mathbb{P}^{3}$ that have points of multiplicity n2n-2 along $r$. This is done in Section 4.
The method of the proof in both cases (a) and (b) is the same, and it is based on a deformation argument. In both cases we construct a flat family $\mathcal{X}\longrightarrow\mathbb{D}$ parametrized by a disk $\mathbb{D}$, such that for $t\neq 0$, the fibre $\mathcal{X}_{t}$ over $t\in\mathbb{D}$ is isomorphic to $\mathbb{P}^{3}$, whereas the central fibre $\mathcal{X}_{0}$ is the union of two smooth irreducible components $A\cup\Theta$, that intersect transversally along a surface $E$. The starting surface $S$ for which we consider Problem 1.1 in the two cases (a) and (b) has a minimal resolution $f:S^{\prime}\longrightarrow S$, with a curve $C$ mapped by $f$ to the singular locus of $S$. What we do is to realize $S^{\prime}$ as a surface $S_{A}$ in the threefold $A$ in such a way that $S_{A}\cap E=C$. Then we construct a surface $S_{\Theta}\subset\Theta$ such that $S_{\Theta}\cap E=C$ so that $S_{0}=S_{A}\cup S_{\Theta}$ is a Cartier divisor in $\mathcal{X}_{0}$. Finally we can do this so that $S_{\Theta}$ has $\delta$ nodes off $E$, that impose independent conditions to the surfaces in the linear system $|S_{0}|$ on $\mathcal{X}_{0}$. If this is the case then we can apply [4, Thm. 4.6] that ensures that $S_{0}$ can be deformed to a surface $S_{t}\subset\mathcal{X}_{t}\cong\mathbb{P}^{3}$ of degree $n$ with $\delta$ nodes that give independent conditions to surfaces of degree $n$. From this it follows that $S$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$.
We notice that the strength of this method consists in studying deformations of a surface $S\subset\mathbb{P}^{3}$ corresponding to a (presumably) singular point of a Severi variety in terms of deformations of a surface $S_{A}\cup S_{\Theta}$ corresponding to a smooth point in a relative Severi variety.
Aknowledgements: The authors are members of GNSAGA of INdAM. In particular, the second author acknowledges funding from the GNSAGA of INdAM and the European Union - NextGenerationEU under the National Recovery and Resilience Plan (PNRR) - Mission 4 Education and research - Component 2 From research to business - Investment 1.1, Prin 2022 “Geometry of algebraic structures: moduli, invariants, deformations”, DD N. 104, 2/2/2022, proposal code 2022BTA242 - CUP J53D23003720006.
The authors also thank Fabrizio Catanese, Barbara Fantechi and Rita Pardini for useful discussions on related problems to the subject of this paper.
2. Nodal surfaces with general plane curve section
Before considering Problem 1.1, in this section we will consider the following related problem:
Problem 2.1.
Let $V$ be an irreducible component of a Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$. Let $S]\in V$ be general and let $C$ be a general plane curve section of the surface $S$. When does it occur that $C$ is a general plane curve of degree $n$?
Let $V$ as above be such that the answer to Problem 2.1 is affermative. Then we will say that $V$ is a good component of the Severi variety.
Proposition 2.2.
Let $V$ be a good component of a Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $n\geqslant 2$. Then
(1) $\dim(V)\geqslant\frac{n(n+3)}{2}.$
Proof.
Let $\mathcal{H}_{n}$ be the Hilbert scheme of plane curves of degree $n$ in $\mathbb{P}^{3}$. One has that $\mathcal{H}_{n}$ is irreducible and $\dim(\mathcal{H}_{n})=\frac{n(n+3)}{2}+3$.
Consider the incidence variety
(2) $\mathbb{I}=\{([S],[C])\in V\times\mathcal{H}_{n}:C\subset S\}.$
One has that $\mathbb{I}$ is irreducible of dimension $\dim(\mathbb{I})=\dim(V)+3$. On the other hand, since $V$ is a good component, the projection $\pi:\mathbb{I}\longrightarrow\mathcal{H}_{n}$ is dominant. Hence we have
$\dim(V)+3=\dim(\mathbb{I})\geqslant\mathcal{H}_{n}=\frac{n(n+3)}{2}+3,$
whence the assertion follows. ∎
Recall that the tangent space to the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ at a point $[S]$ can be identified with $H^{0}(S,\mathcal{O}_{S}(n)\otimes\mathcal{I}_{N})$, where $N$ is the reduced scheme of lenght $\delta$ of the nodes of $S$. Recall that an irreducible component $V$ of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ is said to be regular if for the general point $S]\in V$ one has $h^{1}(S,\mathcal{O}_{S}(n)\otimes\mathcal{I}_{N})=h^{1}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n)\otimes\mathcal{I}_{N})=0$. This means that the points in $N$ give $\delta$ independent conditions to the surfaces in $|\mathcal{O}_{\mathbb{P}^{3}}(n)|$. If $V$ is regular, then $V$ is reduced and
$\dim(V)=h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n))-1-\delta={{n+3}\choose 3}-1-\delta.$
We will say that $V$ is very regular if for the general $S]\in V$ one has $h^{1}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})=0$, with $N$ as above the set of nodes of $S$. Of course, if $V$ is very regular, it is also regular. The following proposition is due to Severi in [11, §1]:
Proposition 2.3.
Let $V$ be a very regular irreducible component $V$ of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $n\geqslant 3$. Then
(3) $\delta\leqslant{{n+2}\choose 3}-4=:\delta_{0}(n).$
Proof.
Let $S]\in V$ be general and let $N$ be its set of double points. Under the hypothesis, one has
$h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})=h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1))-\delta={{n+2}\choose 3}-\delta.$
On the other hand, $h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})\geqslant 4$, because in $H^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})$ there are the first polars of $S$ that are linearly independent because $S$ is not a cone. The assertion follows. ∎
Proposition 2.4.
Let $V$ be a regular irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$. Then $V$ is good if and only if it is very regular.
Proof.
This result is trivially verified for $n=1$. So we may assume $n\geqslant 2$.
Consider again the incidence variety $\mathbb{I}$ as in (2) and the projection $\pi:\mathbb{I}\longrightarrow\mathcal{H}_{n}$. Let $([S],[C])\in\mathbb{I}$ be a general point, and let $N$ be, as usual, the set of $\delta$ nodes of $S$. The tangent space to the fibre of $\pi$ passing through $([S],[C])$ at $([S],[C])$ is
$H^{0}(S,\mathcal{O}_{S}(n)\otimes\mathcal{I}_{N}\otimes\mathcal{I}_{C})=H^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N})=H^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N}).$
Hence, the image of $\pi:\mathbb{I}\longrightarrow\mathcal{H}_{n}$ has dimension
$\dim({\rm Im}(\pi))=\dim(\mathbb{I})-h^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N})=$
$=\dim(V)+3-h^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N})=$
$={{n+3}\choose 3}-\delta+2-h^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N}).$
Now $V$ is good if and only if
$\frac{n(n+3)}{2}+3=\dim(\mathcal{H}_{n})=\dim({\rm Im}(\pi))={{n+3}\choose 3}-\delta+2-h^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N})$
i.e., if and only if
$h^{0}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})=h^{0}(S,\mathcal{O}_{S}(n-1)\otimes\mathcal{I}_{N})=$
$={{n+3}\choose 3}-\delta-1-\frac{n(n+3)}{2}={{n+2}\choose 3}-\delta$
hence if and only if $N$ imposes independent conditions to $|\mathcal{O}_{\mathbb{P}^{3}}(n-1)|$, so if and only if $V$ is very regular. ∎
Remark 2.5.
The regularity hypothesis in Proposition 2.4 is verified for every irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ if either $n\leqslant 7$ or $n\geqslant 8$ and $\delta\leqslant 4n-5$ (see [9, Thm. 1.1]).
In view of the applications of the above results, we record the following proposition.
Proposition 2.6.
Every non-empty irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ is very regular, hence good, if $n\leqslant 5$.
Proof.
The result follows by Proposition 2.4 and by [5, Cor. 1.3]
Remark 2.7.
(a) We recall that the maximal number of nodes of a reduced surface of degree $n\leqslant 5$ in $\mathbb{P}^{3}$ is $4$ for a cubic surface, $16$ for a quartic surface, and $31$ for a quintic surface (see, as a general reference for nodal surfaces in $\mathbb{P}^{3}$, Lab’s thesis [10]). Four nodal cubic surfaces are known as Cayley cubics, $16$-nodal quartic surfaces are known as Kummer surfaces and $31$-nodal quintic surfaces are known as Togliatti surfaces. The Severi varieties $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(3)|}_{4}$ and $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(4)|}_{16}$ are also known to be irreducible.
Notice that for the variety of Kummer surfaces $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(4)|}_{16}$ and for the variety of Togliatti surfaces $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(5)|}_{31}$ the equality holds in (3).
(b) The maximum number of nodes of a sextic surface in $\mathbb{P}^{3}$ is 65 (see, e.g., [10]). A component of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{65}$ is regular by Remark 2.5, but is not very regular, because for it the bound in (3) fails. It is a problem to see if there is a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{\delta}$ for $\delta\leqslant\delta_{0}(6)=52$. More generally it is not known which is the maximal value of $\delta$ such that there exists a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $n\geqslant 6$. In particular, do there exist, for $n\geqslant 4$, very regular components of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta_{0}(n)}$, where $\delta_{0}(n)$ is defined in (3)?
(c) In [4, Thm. 4.6] we proved that there is an irreducible, regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, for any $\delta\leqslant{{n-1}\choose 2}$. It is easy to check that in fact we prove there more than stated, since we prove that there is an irreducible, very regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, for any $\delta\leqslant{{n-1}\choose 2}$.
3. Deformations of surfaces with an ordinary singularity
In this section we consider Problem 1.1 for a reduced surface $S$ of degree $n$ in $\mathbb{P}^{3}$ with a unique ordinary singularity in $p$ of multiplicity $m$. Recall that $S$ has a point of multiplicity $m$ at $p$ if, given the minimal desingularization $f:S^{\prime}\longrightarrow S$ of $S$ at $p$, then, the exceptional divisor of $S^{\prime}$ over $p$ is a plane curve $C$ of degree $m$. If the curve $C$ is smooth, then one says that the singularity of $S$ at $p$ is ordinary. If $C$ is a general plane curve of degree $m$ we will say that $p$ is a general ordinary multiple point. If $C$ is reduced and nodal, we will say that $p$ is a quasi–ordinary multiple point.
We note that the family $\mathcal{T}_{n,m}$ of surfaces of degree $n$ in $\mathbb{P}^{3}$ with a point of multiplicity $m$ is irreducible of dimension
$t_{n,m}={{n+3}\choose 3}-{{m+2}\choose 3}+2=\dim(|\mathcal{O}_{\mathbb{P}^{3}}(n)|)-\delta_{0}(m)-1$
(where $\delta_{0}(m)$ is defined in (3)) and the general point of $\mathcal{T}_{n,m}$ corresponds to a surface with a general ordinary singularity of multiplicity $m$.
In order to attack Problem 1.1 in the case under consideration, we will use a deformation argument that we are going to describe now.
3.1. The deformation
Let $\mathcal{X}^{\prime}=\mathbb{P}^{3}\times\mathbb{D}\longrightarrow\mathbb{D}$ be a trivial family parametrized by the disc $\mathbb{D}=\{t\in\mathbb{C}|\,|t|<1\}$, with fibres $\mathcal{X}^{\prime}_{t}\simeq\mathbb{P}^{3}$ over $t\in\mathbb{D}$.
Next we blow-up the point $p\in\mathcal{X}^{\prime}_{0}$. Denote by $\mathcal{X}\longrightarrow\mathcal{X}^{\prime}$ the blowing-up of $\mathcal{X}^{\prime}$ at $p\in\mathcal{X}^{\prime}_{0}$. Then $\mathcal{X}$ has fibre $\mathcal{X}_{t}\simeq\mathbb{P}^{3}$ if $t\neq 0$ and central fibre is
$\mathcal{X}_{0}=A\cup\Theta,$
where $A$ is the blowing-up $p:A\longrightarrow\mathbb{P}^{3}$ of $\mathbb{P}^{3}$ at $p$ and $\Theta\cong\mathbb{P}^{3}$ is the exceptional divisor, intersecting transversally $A$ along the plane $E=A\cap\Theta$.
Let now $S\subset\mathcal{X}^{\prime}_{0}\cong\mathbb{P}^{3}$ be a degree $n$ surface with an ordinary or quasi–ordinary point of multiplicity $m$ at $p$ and at most nodes as further singularities. We denote by $H$ the Cartier divisor on $\mathcal{X}$ which is pull-back of a general plane divisor of $\mathbb{P}^{3}$ with respect to the projection morphism $\mathcal{X}\longrightarrow\mathbb{P}^{3}$. The strict transform $S_{A}$ of $S\subset\mathcal{X}^{\prime}_{0}$ on $\mathcal{X}$ is nothing but the minimal resolution $S^{\prime}$ of $S$ at $p$, it sits on the component $A$ of $\mathcal{X}_{0}$ and it belongs to the linear system $|\mathcal{O}_{A}(nH-mE)|$. Under our hypotheses on $S$, the intersection $C=S_{A}\cap E$ is a reduced plane curve of degree $m$ in $E\cong\mathbb{P}^{2}$ with at most nodes.
Let now $S_{0}:=S_{A}\cup S_{\Theta}\in|\mathcal{O}_{\mathcal{X}_{0}}(nH-m\Theta)|$, be a Cartier divisor with $S_{\Theta}\subset\Theta$. Then, observing that $\Theta\cong\mathbb{P}^{3}$ and $\Theta|_{\Theta}:=H_{\Theta}$ is a plane, we have that $S_{\Theta}\in|mH_{\Theta}|$ is a surface of degree $m$ in $\Theta\cong\mathbb{P}^{3}$, that intersects the plane $E$ along the curve $C$. We require that $S_{\Theta}$ is smooth along $C$, so that $S_{\Theta}$ has at most isolated singularities off $E$.
Note that $\mathcal{O}_{\mathcal{X}}(nH-m\Theta)$ is a line bundle on $\mathcal{X}$ such that its restriction to $\mathcal{X}_{t}$, for $t\neq 0$, is $\mathcal{O}_{\mathcal{X}_{t}}(nH)\cong\mathcal{O}_{\mathbb{P}^{3}}(n)$.
3.2. Nodal deformations
In what follows we will assume that $S_{\Theta}$ is nodal and we will be interested in nodal deformations of a surface of the type of $S_{0}$ in $\mathcal{X}$.
Proposition 3.1.
In the above set–up, suppose that $C$ has $\delta_{C}$ nodes, that $S_{A}$ has $\delta_{A}$ nodes and that $S_{\Theta}$ has $\delta_{\Theta}$ nodes all off $E$. Let us denote by $N$ the set of these $\delta:=\delta_{A}+\delta_{C}+\delta_{\Theta}$ nodes. Suppose that $h^{1}(\mathcal{X}_{0},\mathcal{O}_{\mathcal{X}_{0}}(nH-m\Theta)\otimes I_{N})=0$, that is equivalent to say that the points in $N$ give independent conditions to surfaces in $|\mathcal{O}_{\mathcal{X}_{0}}(nH-m\Theta)|$. Then $S_{0}$ can be deformed to a $\delta$–nodal surface $S_{t}\subset\mathcal{X}_{t}\cong\mathbb{P}^{3}$ of degree $n$, for $t\neq 0$ and the point corresponding to $S$ sits in the closure of a regular component of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$.
Proof.
The proposition follows as an immediate application of [4, Thm. 3.13]. ∎
3.3. Applications
3.3.1. The case of general ordinary multiple point
Our first application is to the case $p$ is a general ordinary multiple point of multiplicity $m$ for $S$.
Theorem 3.2.
Let $\delta$ be such that there is a very regular irreducible component $V$ of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(m)|}_{\delta}$. Then the variety $\mathcal{T}_{n,m}$ sits in the Zariski closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$. In other words, the general degree $n$ surface $S$ with a general ordinary singularity of multiplicity $m$ is the limit of surfaces of degree $n$ with $\delta$ nodes tending to the multiplicity $m$ point.
Proof.
Let $S]\in\mathcal{T}_{n,m}$ be a general point. The corresponding surface $S\subset\mathbb{P}^{3}$ has a unique multiple point which is a general ordinary singularity of multiplicity $m$. If we perform the construction we made in Section 3.1, the curve $C$ cut out by the strict transform $S_{A}$ of $S$ on $E=A\cap\Theta$ is a general plane curve of degree $m$. Now, if $\delta$ and $V$ are as in the statement, we can find $S_{\Theta}\in|mH_{\Theta}|=|\mathcal{O}_{\mathbb{P}^{3}}(m)|$ cutting $C$ on $E$, with $\delta$ nodes lying in $V$ because $V$ is good by Proposition 2.4. Moreover, since $V$ is also regular, the nodes of $S_{\Theta}$ impose independent conditions to the surfaces in $|\mathcal{O}_{\mathbb{P}^{3}}(m)|$. Hence the assertion follows by Proposition 3.1. ∎
In order to prove a corollary of this theorem, we need some preliminary results.
Proposition 3.3.
Let $(S,p)$ be a germ of surface singularity in $(\mathbb{C}^{3},\bf 0)$, with $p$ an isolated point of multiplicity 3 with tangent cone $\Gamma$ with vertex $p$ over a smooth curve $C$ of degree 3 in $\mathbb{P}^{2}$. Then $(S,p)$ is analitically equivalent to $(\Gamma,p)$.
Proof.
Suppose that $\Gamma$ has equation $f(x,y,z)=0$, where $f(x,y,z)=0$ is a homogeneous polynomial of degree 3, that defines $C$ in $\mathbb{P}^{2}$. By the finite determinacy theorem [7, Thm. 2.23], $(S,p)$ is analitically isomorphic to $(\Gamma,p)$ if
(4) $\mathfrak{m}^{4}\subseteq\mathfrak{m}^{2}\cdot\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)$
where $\mathfrak{m}$ is the maximal ideal (x,y,z) in $\mathbb{C}[x,y,z]$. Let us prove that (4) holds with in fact the equality. Indeed, consider the multiplication map
$\mu:\mathbb{C}[x,y,z]_{2}\otimes\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)\longrightarrow\mathbb{C}[x,y,z]_{4},$
where $\mathbb{C}[x,y,z]_{d}$ is the vector space of homogeneous polynomials of degree $d$. Since $C$ is smooth, $\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)$ is a regular sequence, hence the kernel of $\mu$ has exactly dimension 3, being generated by the Koszul syzygies of $\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}$. Thus
$\dim({\rm Im}(\mu))=3\dim(\mathbb{C}[x,y,z]_{2})-3=15=\dim(\mathbb{C}[x,y,z]_{4})$
that implies the assertion.∎
Lemma 3.4.
Let $S$ be a cubic surface in $\mathbb{P}^{3}$ with an isolated singularity at $p$. Then the natural map
$H^{0}(S,N_{S|\mathbb{P}^{3}})\longrightarrow H^{0}(T^{1}_{S,p})$
is surjective. In particular all local deformations of the singularity $(S,p)$ can be realised by deforming $S$ as a cubic in $\mathbb{P}^{3}$.
Proof.
We have the long exact sequence of sheaves
$0\longrightarrow T_{S}\longrightarrow T_{\mathbb{P}^{3}|_{S}}\longrightarrow N_{S|\mathbb{P}^{3}}\longrightarrow T^{1}_{S,p}\longrightarrow 0.$
We denote by $N^{\prime}_{S|\mathbb{P}^{3}}$ the kernel of the map $N_{S|\mathbb{P}^{3}}\longrightarrow T^{1}_{S,p}$, so that we have the two short exact sequences
(5) $0\longrightarrow T_{S}\longrightarrow T_{\mathbb{P}^{3}|_{S}}\longrightarrow N^{\prime}_{S|\mathbb{P}^{3}}\longrightarrow 0$
and
(6) $0\longrightarrow N^{\prime}_{S|\mathbb{P}^{3}}\longrightarrow N_{S|\mathbb{P}^{3}}\longrightarrow T^{1}_{S,p}\longrightarrow 0.$
We claim that $h^{2}(S,T_{S})=0$. Indeed by Serre duality $h^{2}(S,T_{S})=h^{0}(S,\Omega_{S}(-1))$. So it suffices to prove that $h^{0}(S,\Omega_{S})=0$.
We have the exact sequence
$0\longrightarrow\mathcal{O}_{S}(-3)\longrightarrow\Omega_{\mathbb{P}^{3}|_{S}}\longrightarrow\Omega_{S}\longrightarrow 0.$
By the vanishing of $h^{1}(S,\mathcal{O}_{S}(-3))=h^{1}(S,\mathcal{O}_{S}(2))=0$, the assertion will follow by proving that $h^{0}(S,\Omega_{\mathbb{P}^{3}|_{S}})=0$. We look at the exact sequence
(7) $0\longrightarrow\Omega_{\mathbb{P}^{3}}(-3)\longrightarrow\Omega_{\mathbb{P}^{3}}\longrightarrow\Omega_{\mathbb{P}^{3}|_{S}}\longrightarrow 0.$
We have clearly $h^{0}(\mathbb{P}^{3},\Omega_{\mathbb{P}^{3}})=0$ and also $h^{1}(\mathbb{P}^{3},\Omega_{\mathbb{P}^{3}}(-3))=0$ by Bott’s formula. Hence $h^{0}(S,\Omega_{\mathbb{P}^{3}|_{S}})=0$.
We claim next that $h^{1}(S,T_{\mathbb{P}^{3}|_{S}})=0$. In fact we have the Euler sequence
$0\longrightarrow\mathcal{O}_{S}\longrightarrow\mathcal{O}_{S}(1)^{\oplus 4}\longrightarrow T_{\mathbb{P}^{3}|_{S}}\longrightarrow 0.$
Since $h^{1}(S,\mathcal{O}_{S}(1))=0$ and $h^{2}(S,\mathcal{O}_{S})=h^{0}(S,\omega_{S})=h^{0}(S,\mathcal{O}_{S}(-1))=0$, the assertion follows.
Then, since $h^{1}(S,T_{\mathbb{P}^{3}|_{S}})=h^{2}(S,T_{S})=0$, from (5) follows that $h^{1}(S,N^{\prime}_{S|\mathbb{P}^{3}})=0$. Then the assertion follows from (6). ∎
By applying Proposition 2.6, Remark 2.7, (c), Proposition 3.3 and Lemma 3.4, we have:
Corollary 3.5.
One has:
  • (i)
    $\mathcal{T}_{n,3}$ sits in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ as a subvariety of codimension 3, but not in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ with $\delta>4$. More precisley, an isolated ordinary triple point of a surface in $\mathbb{P}^{3}$ is never a limit of more than four nodes;
  • (ii)
    $\mathcal{T}_{n,4}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$, as a codimension $1$ subvariety;
  • (iii)
    $\mathcal{T}_{n,5}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$, as a codimension $1$ subvariety;
  • (iv)
    for any $m\geqslant 5$, $\mathcal{T}_{n,m}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{{m-1}\choose 2}$.
Note that Proposition 3.3 and Lemma 3.4 are used only for (i).
Remark 3.6.
Concerning point (i) of Corollary 3.5, notice that $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ is clearly irreducible and regular. By [9, Thm 1.1], any irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$ [resp., of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$] is regular as soon as $n\geqslant 6$ [resp., as soon as $n\geqslant 9$].
3.3.2. The case of a (not necessarily general) ordinary singularity
We have:
Theorem 3.7.
Let $S$ be a surface of degree $n$ in $\mathbb{P}^{3}$ with a unique ordinary singularity $p$ of multiplicity $m$. Consider the minimal desingularization $f:S^{\prime}\longrightarrow S$ of $S$ with the exceptional divisor of $S^{\prime}$ over $p$ a plane curve $C$ of degree $m$. Let $\delta$ be such that there is a surface $X$ of degree $m$ in $\mathbb{P}^{3}$ having $C$ as a plane section and belonging to a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(m)|}_{\delta}$. Then the point corresponding to $S$ belongs to the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$.
Proof.
Again this is an immediate consequence of Proposition 3.1. ∎
Remark 3.8.
Let $C$ be a smooth plane curve of degree $m$. Theorem 3.7 poses the problem of determining the maximum $\delta(C)$ of $\delta$ such that there is a surface $X$ of degree $m$ in $\mathbb{P}^{3}$ having $C$ as a plane section and belonging to a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(m)|}_{\delta}$.
If $C$ is general, one has $\delta(C)\leqslant\delta_{0}(m)$. In fact, by Proposition 2.4, a regular and good component of the Severi variety is also very regular, and therefore the assertion follows by Proposition 2.3. As in Remark 2.7, (b), one can ask whether $\delta(C)=\delta_{0}(m)$ if $C$ is general. Suppose that this is the case. Then, by the proof of Theorem 3.7, it would follow that $\mathcal{T}_{n,m}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta_{0}(m)}$ as a divisor.
Can the invariant $\delta(C)$ vary with $C$?
3.4. An interesting example of a quasi–ordinary multiple point
In this section, we want to show that the answer to Problem 1.1 certainly depends on whether the point $p\in S$ ($S$ as usual of degree $n$ with no other singularity) is ordinary or quasi–ordinary. We make this by considering the example of a triple point.
If the triple point $p\in S$ is ordinary, then, as we saw, the point corresponding to $S$ is the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ and in general it does not sit in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ with $\delta>4$ (see Corollary 3.5, (i)). We will now show that there are quasi–ordinary triple points $p\in S$ such that $S$ is the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{7}$.
Lemma 3.9.
Let $X\subset\mathbb{P}^{3}$ be a 4–nodal cubic. There is a plane section of $X$, not containing any of the four nodes, that consists of three lines not in a pencil.
The seven points consisting of the four nodes and of the three vertices of this triangle give independent conditions to cubics in $\mathbb{P}^{3}$.
Proof.
It is well known that $X$ is the image of the plane via the rational map determined by the linear system of cubic curves passing through the vertices $P_{1},\ldots,P_{6}$ of a quadrilateral.
Figure 1.
In Figure 1 the lines $L_{N_{1}},\ldots,L_{N_{4}}$ of the quadrilateral are depicted in black. They are mapped to the four nodes $N_{1},\ldots,N_{4}$ respectively.
Look at the three blue lines $L_{1},L_{2},L_{3}$ in Figure 1. Each of them is mapped to a line and the union of the three lines forms a cubic through $P_{1},\ldots,P_{6}$. This proves the first assertion of the lemma.
As for the second assertion, we consider the linear system $\mathcal{L}$ of cubics passing through the four nodes of $S$. This has dimension $15$ because the nodes impose independent conditions to the cubics. Consider the restriction of the linear system $\mathcal{L}$ to the plane containing $L_{1},L_{2},L_{3}$. The kernel of this restriction is isomorphic to the linear system of quadrics passing through the four nodes, that has dimension 5. So the restriction linear system has dimension 9 hence it is the complete linear system of cubics in the plane, and this implies the assertion. ∎
Proposition 3.10.
Consider a surface $S\subset\mathbb{P}^{3}$ of degree $n\geqslant 4$ with, as its only singularity, a quasi–ordinary triple point $p$, whose minimal resolution $f:S^{\prime}\longrightarrow S$ is such that the exceptional divisor $C$ over $p$ is a plane cubic with three nodes, i.e., it is a plane cubic consisting of three lines not in a pencil. Then $S$ is in the closure of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{7}$.
Proof.
Let us go back to the deformation set–up as in Section 3.1. We have $S_{A}=S^{\prime}$, and we may assume that $S_{\Theta}$ is the 4–nodal cubic surface having as plane section the triangle $C$, as shown in Lemma 3.9. Then the assertion follows from Proposition 3.1. ∎
Remark 3.11.
Again by the results in [9], any irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{7}$ is regular.
4. Some rational surfaces in $\mathbb{P}^{3}$
In this section we will introduce and study some irreducible rational surfaces $S$ of degree $n\geqslant 3$ in $\mathbb{P}^{3}$ that have a line $R$ of points of multiplicity n2n-2. These surfaces will be used in Section 5 when we will consider Problem 1.1 for surfaces with a double line.
4.1. A remark on hyperelliptic curves
Let $n\geqslant 3$ be an integer. Let us consider the linear system $\mathcal{C}$ of dimension 3n43n-4 of plane curves of degree n1n-1 with a fixed point $q\in\mathbb{P}^{2}$ of multiplicity at least n3n-3. For $n=3$ we are just considering the complete linear system of conics and a fixed point $q$ in the plane. There is a dense open subset $U$ of $\mathcal{C}$ such that all points in $U$ correspond to curves that have an ordinary singularity of multiplicity n3n-3 at $q$ and are elsewhere smooth, so that the normalization of these curves have genus $p:=n-3\geqslant 0$ and have the $g^{1}_{2}$ cut out by the lines through $q$. If $G$ is the group of projectivities that fixes $q$, we have an obvious morphism $f:U/G\longrightarrow\mathcal{M}^{1}_{p,2}$ where $\mathcal{M}^{1}_{p,2}$ is the moduli space of pairs $(C,\mathfrak{h})$ with $C\in\mathcal{M}_{p}$ a curve of genus $p$ and $\mathfrak{h}$ a $g^{1}_{2}$ on $C$. Note that $\mathcal{M}^{1}_{p,2}$ coincides with the locus in $\mathcal{M}_{p}$ of hyperelliptic curves, if $p\geqslant 2$, whereas $\mathcal{M}^{1}_{1,2}$ is the relative Picard bundle of degree 2 line bundles on $\mathcal{M}_{1}$, and, finally, $\mathcal{M}^{1}_{0,2,}\simeq{\mathbb{P}^{2}}^{*}$ is the set of $g^{1}_{2}$ on $\mathbb{P}^{1}$. In particular, the dimension of $\mathcal{M}^{1}_{p,2}$ is 2p12p-1 if $p\geqslant 2$, and it is $2$ if $p=0,1$.
Proposition 4.1.
The map $f:U/G\longrightarrow\mathcal{M}^{1}_{p,2}$ is dominant.
Proof.
Take $(C,\mathfrak{h})$ general in $\mathcal{M}^{1}_{p,2}$. Consider a general, non special, $g^{1}_{n-2}$ on $C$, that we denote by $\mathfrak{k}$. The minimal sum of $\mathfrak{h}$ and $\mathfrak{k}$ has dimension $3$ and maps $C$ isomorphically to a curve of degree $n$ and type (2,n2)(2,n-2) on a smooth quadric in $\mathbb{P}^{3}$. Let $x\in C$ be a general point, and project $C$ down from $x$ to a plane $\mathbb{P}^{2}$. The image $\Gamma$ of $C$ is a curve of degree n1n-1 that has a point of multiplicity n3n-3 as wanted. ∎
Remark 4.2.
Assume $n\geqslant 5$, i.e., $p\geqslant 2$. Then the general fibre of $f$ has dimension
$\dim(U/G)-\dim(\mathcal{M}^{1}_{p,2})=(3n-10)-(2p-1)=n-3=p$
which means that given a general hyperelliptic curve $C$ of genus p=n3p=n-3, it can be realized as a plane curve of degree n1n-1 with a point of multiplicity n3n-3 and no other singularity in $\infty^{p}$ ways, i.e., the degree n1n-1 embedding line bundle in the plane varies in the whole ${\rm Pic}^{n-1}(C)$. Of course the same conclusion is trivially true for $n=3$, and it also holds for $n=4$. In these two cases, the general fibre of $f$ has dimension $\dim(U/G)-\dim(\mathcal{M}^{1}_{p,2})=0.$
4.2. Certain linear systems of plane curves
Let us consider a plane curve $\Gamma$ of degree $n-1\geqslant 2$ with an ordinary point $q$ of multiplicity n3n-3 and no other singularity, with normalization $C\longrightarrow\Gamma$ a curve of genus p=n3p=n-3 and let us consider the linear series $\mathfrak{h}=g^{1}_{2}$ on $C$ cut out by the lines through $q$. Suppose that $(C,\mathfrak{h})$ is a general point in $\mathcal{M}^{1}_{p,2}$. The ramification points of $\mathfrak{h}$ are 2n42n-4. Hence there are 2n42n-4 distinct lines through $q$ that are tangent to $\Gamma$ in points that, by generality, we can assume to be off $q$.
Let us write $3n-4=2\ell+\epsilon$, with $0\leqslant\epsilon\leqslant 1$, and notice that $4n-8\geqslant 2\ell$. Indeed, the assertion is trivial for $n=3$ and if $n\geqslant 4$ then $4n-8\geqslant 3n-4\geqslant 2\ell$. Thus, we can fix $\ell\leqslant 2n-4$ of the 2n42n-4 tangent lines to $\Gamma$ passing through $q$ and let us denote them by $r_{1},\ldots,r_{\ell}$. Let us denote by $q_{1},\ldots,q_{\ell}$ the tangency points of $r_{1},\ldots,r_{\ell}$ with $\Gamma$ off $q$. Let us also denote by $q^{\prime}_{i}$ the infinitely near point to $q_{i}$ along $r_{i}$, for $i=1,\ldots,\ell$, and notice that the $2\ell$ points $q_{i},q^{\prime}_{i}$, for $i=1,\ldots,\ell$, belong to $\Gamma$. If $\epsilon=1$ let us fix also a general point $\bar{q}$ on $\Gamma$.
Consider now the linear system $\mathcal{L}$ of plane curves of degree $n$ that have at $q$ a point of multiplicity at least n2n-2 and pass through the points $q_{i},q^{\prime}_{i}$, for $i=1,\ldots,\ell$, and through $\bar{q}$ if $\epsilon=1$.
We want to study the system $\mathcal{L}$ and the image of the plane via the rational map determined by $\mathcal{L}$. To do this, it is convenient to pass to the blow–up $f:X\longrightarrow\mathbb{P}^{2}$ of the plane at the point $q$, at the points $q_{i}$ and then at the infinitely near points $q^{\prime}_{i}$, for $i=1,\ldots,\ell$, and at $\bar{q}$ if $\epsilon=1$. We let $\mathcal{L}^{\prime}$ be the line bundle on $X$ corresponding to the strict transform of $\mathcal{L}$ on $X$. Let $R$ be the pull–back on $X$ of a general line of $\mathbb{P}^{2}$, let $L$ be the strict transform on $X$ of a general line of $\mathbb{P}^{2}$ through $q$ and we denote by $C$ the strict transform on $X$ of the curve $\Gamma$, that is the normalization of $\Gamma$. Then we have $\mathcal{O}_{C}(L)\sim\mathfrak{h}$ and, as we observed, $\mathcal{O}_{C}(R)$ is a general line bundle on $C$ of degree n1n-1.
Consider the exact sequence
(8) $0\longrightarrow\mathcal{L}^{\prime}(-C)\longrightarrow\mathcal{L}^{\prime}\longrightarrow\mathcal{L}^{\prime}_{|C}\longrightarrow 0.$
We notice that $\mathcal{L}^{\prime}(-C)\cong\mathcal{O}_{X}(L)$, and therefore $h^{1}(X,\mathcal{L}^{\prime}(-C))=0$. This implies that $|\mathcal{L}^{\prime}|$ cuts out on $C$ a complete linear series $\mathfrak{g}$. The degree of $\mathfrak{g}$ is
n(n1)(n2)(n3)(3n4)=n2=p+1n(n-1)-(n-2)(n-3)-(3n-4)=n-2=p+1
and more precisely one has that
$\mathfrak{g}\sim(n-2)\mathfrak{h}+2R_{|C}-\ell\mathfrak{h}-\epsilon\bar{q}$
and, as $\mathcal{O}_{C}(R)$ varies in ${\rm Pic}^{n-1}(C)$, we see that $\mathfrak{g}$ varies in the whole of ${\rm Pic}^{n-2}(C)$, so that with an appropriate choice of $\mathcal{O}_{C}(R)$ we may assume that $\mathfrak{g}$ is any given general $g^{1}_{n-2}$ on $C$.
4.3. The image surface
Looking at the sequence (8), under the above generality assumptions on $(C,\mathfrak{h})$ and on the series $\mathfrak{g}$, we see that $\dim(|\mathcal{L}^{\prime}|)=3$, hence also $\dim(\mathcal{L})=3$.
Lemma 4.3.
Under the above generality assumptions on $(C,\mathfrak{h})$ and on the series $\mathfrak{g}$, the linear system $|\mathcal{L}^{\prime}|$ has no base points.
Proof.
The linear system $|\mathcal{L}^{\prime}|$ contains the subsystem consisting of the fixed curve $C$ plus the base point free pencil $|L|$. So the only base points of $|\mathcal{L}^{\prime}|$ may occur on $C$. However, as we saw, $|\mathcal{L}^{\prime}|$ cuts out on $C$ the complete, base point free linear series $\mathfrak{g}$, so $|\mathcal{L}^{\prime}|$ has no base points. ∎
Proposition 4.4.
Under the above generality assumptions on $(C,\mathfrak{h})$ and on the series $\mathfrak{g}$, the linear system $|\mathcal{L}^{\prime}|$ determines a birational morphism of $X$ to a surface $\Sigma$ in $\mathbb{P}^{3}$ of degree $n$ with a line $r$, the image of the curve $C$, of multiplicity n2n-2. Via this morphism, the linear system $|L|$ is mapped to the linear system of conics cut out on $\Sigma$ by the planes through the line $r$.
Proof.
Let $\varphi:X\longrightarrow\Sigma$ be the morphism determined by $|\mathcal{L}^{\prime}|$. Since $|\mathcal{L}^{\prime}|$ contains the subsystem consisting of the fixed curve $C$ plus the base point free pencil $|L|$, to see that $\varphi$ is birational onto its image, it suffices to prove that given the general curve $L\in|L|$, $|\mathcal{L}^{\prime}|$ cuts out on it a complete linear series of degree 2. To see this, look at the sequence
(9) $0\longrightarrow\mathcal{L}^{\prime}(-L)\longrightarrow\mathcal{L}^{\prime}\longrightarrow\mathcal{L}^{\prime}_{|L}\longrightarrow 0,$
with $\mathcal{L}^{\prime}(-L)=\mathcal{O}_{X}(C)$. Then the assertion will be proved if we show that $h^{1}(\mathcal{O}_{X}(C))=0$ or equivalently, by Riemann–Roch theorem, that $h^{0}(\mathcal{O}_{X}(C))=1$. Look at the sequence
$0\longrightarrow\mathcal{O}_{X}\longrightarrow\mathcal{O}_{X}(C)\longrightarrow\mathcal{O}_{C}(C)\longrightarrow 0.$
To prove that $h^{0}(\mathcal{O}_{X}(C))=1$ one has to show that $h^{0}(\mathcal{O}_{C}(C))=0$. Note that $C^{2}=n-4$ and more precisely
$\mathcal{O}_{C}(C)\sim(n-3-\ell)\mathfrak{h}+2R_{|C}-\epsilon\bar{q}.$
Using again the generality of $R_{|C}$, this is a general line bundle of degree n4=p1n-4=p-1 on $C$, and therefore it is not effective, as wanted.
The degree of the image surface $\Sigma$ is
$\operatorname{deg}(\Sigma)=(C+L)^{2}=(n-4)+2\cdot 2=n.$
Finally, recalling that $\mathcal{L}^{\prime}_{|C}=\mathfrak{g}$ is a $g^{1}_{n-2}$ on $C$ and that $(C+L)\cdot L=C\cdot L=2$, we see that $C$ is mapped by $\varphi$ to a line $r$ contained in $\Sigma$ with multiplicity n2n-2, and the linear system $|L|$ is mapped by $\varphi$ to the linear system of conics in $\Sigma$ cut out by the planes through $r$. ∎
4.4. Nodes
Finally we show that:
Proposition 4.5.
The surface $\Sigma$ has $2\ell$ nodes as singularities off the line $r$ of multiplicity n2n-2 that lie pairwise on $\ell$ disjoint lines contained in $\Sigma$ and meeting $r$ at $\ell$ distinct points. These nodes of $\Sigma$ impose independent conditions to the linear system of surfaces of degree $n$ that contain $r$ with multiplicity n2n-2.
Proof.
To prove that $\Sigma$ has $2\ell$ nodes off the line $r$ of multiplicity n2n-2, we will prove that on $X$ there are (2)(-2)–curves $D_{1},\ldots,D_{2\ell}$, such that $\mathcal{L}^{\prime}\cdot D_{i}=C\cdot D_{i}=0$, for $i=1,\ldots,2\ell$, and $D_{i}\cdot D_{j}=0$ for $1\leqslant i<j\leqslant 2\ell$. To see this, notice that the blow–up of the points $q_{1},\ldots,q_{\ell}$, produces (1)(-1)–curves $E_{1},\ldots,E_{\ell}$. When we blow–up the infinitely near points $q^{\prime}_{1},\ldots,q^{\prime}_{\ell}$, with $q^{\prime}_{i}\in E_{i}$, for $i=1,\ldots,\ell$, introducing new exceptional divisors $F_{1},\ldots,F_{\ell}$, the proper transforms of $E_{1},\ldots,E_{\ell}$ on $X$ become (2)(-2)–curves $D_{1},\ldots,D_{\ell}$, such that $\mathcal{L}^{\prime}\cdot D_{i}=C\cdot D_{i}=0$ for $i=1,\ldots,\ell$, and $D_{i}\cdot D_{j}=0$ for $1\leqslant i<j\leqslant\ell$. On the other hand, after subsequently blowing up $q$, the proper transforms $D_{\ell+1},\ldots,D_{2\ell}$ of the lines $r_{1},\ldots,r_{\ell}$ become also (2)(-2)–curves, such that $\mathcal{L}^{\prime}\cdot D_{i}=C\cdot D_{i}=0$ for $i=\ell+1,\ldots,2\ell$, and $D_{j}\cdot D_{j}=0$ for $1\leqslant i<j\leqslant 2\ell$.
To see that the $2\ell$ nodes of $\Sigma$ obtained by the contraction of the (2)(-2)–curves $D_{1},\ldots,D_{2\ell}$ are pairwise contained into $\ell$ disjoint lines contained in $\Sigma$ and meeting $r$ at $\ell$ distinct points, observe that the (1)(-1)–curves $F_{1},\ldots,F_{\ell}$ are such that $F_{i}\cdot F_{j}=0$ if $i\neq j$, $F_{i}\cdot D_{i}=F_{i}\cdot D_{i+l}=1$, $F_{i}\cdot D_{j}=0$ for $j\neq i,i+l$ and $F_{i}\cdot\mathcal{L}^{\prime}=C\cdot F_{i}=1.$ Thus their image curves in $\mathbb{P}^{3}$ by the morphism defined by $\mathcal{L}^{\prime}$ are $\ell$ disjoint lines $s_{1},\ldots,s_{\ell}$ with the desired properties.
We observe that, if $\epsilon=1$ the blowing-up at the point $\bar{q}$ produces two further (1)(-1)-curves on $X$, the corresponding exceptional divisor $\bar{E}$ and the proper transform of the line $\bar{r}$ passing through $q$ and $\bar{q}$. One has
$\bar{E}\cdot\bar{r}=\mathcal{L}^{\prime}\cdot\bar{E}=\mathcal{L}^{\prime}\cdot\bar{r}=C\cdot\bar{E}=C\cdot\bar{r}=1.$
This implies that the morphism defined by $\mathcal{L}^{\prime}$ maps isomorphically both $\bar{E}$ and $\bar{r}$ to lines in $\Sigma\subset\mathbb{P}^{3}$, which we denote respectively by $\bar{e}$ and $\bar{s}$ . We claim that $\bar{e}$ and $\bar{s}$ are two distinct lines, generating a plane $\bar{\pi}$ passing through the multiple line $r$ of $\Sigma$ and tangent to $\Sigma$ at the point $\bar{e}\cap\bar{s}$, which is a smooth point for $\Sigma$. Indeed these two lines have to meet at the image of the intersection point of $\bar{E}$ and $\bar{r}$ and they meet the multiple line $r$ of $\Sigma$ at two distinct points, because, by the generality of $\bar{q}$, we may assume that the two intersection points of the line $\bar{r}$ with $C$ off the point $q$ are not in a divisor of $|\mathcal{L}^{\prime}|_{C}|$ (which we know to be a complete general $g^{1}_{n-2}$ on $C$).
To show that $\Sigma$ has no further singular points outside the line $r$, we first observe that $\Sigma$ may have only singular points of multiplicity $2$ off $r$. This is because every plane $\pi$ passing through $r$ intersects $\Sigma$ along $r$, with multiplicity n2n-2, and a conic. If $\pi_{1},\ldots,\pi_{\ell}$ are the planes generated respectively by the line $r$ and the line $s_{i}$, with $i=1,\dots,\ell$, then every $\pi_{i}$ intersects $\Sigma$ along the line $r$, with multiplicity n2n-2 and the line $s_{i}$, with multiplicity $2$, with $i=1,\dots,\ell$.
By Lemma 4.3 and by the exact sequence (9), under our generality assumptions, for any line $l\in|L|$, different from $r_{1},\ldots,r_{\ell}$, and from $\bar{r}$ if $\epsilon=1$, the linear system $|\mathcal{L}^{\prime}|_{l}|$ is a complete, base points free linear system of dimension and degree $2$. Then the morphism defined by $\mathcal{L}^{\prime}$ isomorphically maps $l$ to a smooth conic. This implies that any plane $\pi$ in $\mathbb{P}^{3}$ passing through $r$ other than $\pi_{1},\ldots,\pi_{\ell}$, and than $\bar{\pi}$ if $\epsilon=1$, intersects $\Sigma$ outside $r$ along a smooth conic. This proves that $\Sigma$ is smooth outside $r$ and the $2\ell$ aforementioned nodes.
We are now left to prove the final assertion. For this, we must prove that the (2)(-2)–curves $D_{1},\ldots,D_{2\ell}$ give $2\ell$ independent conditions to the pull–back $\mathcal{M}$ on $X$ via $\varphi$ of the linear system of surfaces of degree $n$ in $\mathbb{P}^{3}$ that contain $r$ with multiplicity n2n-2. One has
$\mathcal{M}\sim n\mathcal{L}^{\prime}-(n-2)C\sim 2\mathcal{L}^{\prime}+(n-2)L$
that corresponds to the linear system of plane curves of degree 3n23n-2 with the point $q$ of multiplicity 3n63n-6, with double points at the points $q_{1},\ldots,q_{\ell}$ and $q^{\prime}_{1},\ldots,q^{\prime}_{\ell}$, and a further double point at $\bar{q}$ if $\epsilon=1$.
Claim 4.6.
The linear system $\mathcal{M}$ has dimension 6n46n-4.
Proof of Claim 4.6.
First we estimate the dimension of the linear system $\mathcal{S}$ of surfaces of degree $n$ in $\mathbb{P}^{3}$ having a line $r$ of multiplicity (at least) n2n-2. Take a general plane $\pi$ containing $r$. The system $\mathcal{S}$ cuts out on $\pi$, off $r$, a linear system of conics. Hence we can split $\pi$ off $\mathcal{S}$ with at most $6$ conditions. Repeating this n2n-2 times, the residual system is the system of all quadrics, that has dimension 9. Hence we have
$\dim(\mathcal{S})\leqslant 6(n-2)+9=6n-3.$
Accordingly, we must have $\dim(\mathcal{M})\leqslant 6n-4$. On the other hand
$\dim(\mathcal{M})\geqslant\frac{(3n-2)(3n+1)}{2}-\frac{(3n-6)(3n-5)}{2}-6\ell-3\epsilon=6n-4,$
hence $\dim(\mathcal{M})=6n-4$, as wanted. ∎
Now subtract from $\mathcal{M}$ the $\ell$ lines $r_{1},\ldots,r_{\ell}$. We are left with a linear system $\mathcal{N}$ and
$\dim(\mathcal{N})\geqslant\dim(\mathcal{M})-\ell=6n-4-\ell.$
The linear system $\mathcal{N}$ corresponds to the linear system of plane curves of degree 3n23n-2-$\ell$ with the point $q$ of multiplicity 3n63n-6-$\ell$, with simple points at the points $q_{1},\ldots,q_{\ell}$ and $q^{\prime}_{1},\ldots,q^{\prime}_{\ell}$, and a double point at $\bar{q}$ if $\epsilon=1$. Now we want to split from $\mathcal{N}$ also the remaining (2)(-2)-curves, and this gives us the linear system $\mathcal{P}$ that corresponds to the linear system of plane curves of degree 3n23n-2-$\ell$ with the point $q$ of multiplicity 3n63n-6-$\ell$, with double points at $q_{1},\ldots,q_{\ell}$ and a double point at $\bar{q}$ if $\epsilon=1$. We have
(10) $\dim(\mathcal{P})\geqslant\dim(\mathcal{N})-\ell\geqslant\dim(\mathcal{M})-2\ell=6n-4-2\ell=3n+\epsilon.$
By the generality assumptions, we may assume that the points $q_{1},\ldots,q_{\ell}$ and $\bar{q}$ (if $\epsilon=1)$ are general points in the plane.
Consider now the surface $Y\subset\mathbb{P}^{r}$, with $r=15n-5\ell-16$, that is the image of the plane via the linear system of curves of degree 3n23n-2-$\ell$ with the point $q$ of multiplicity 3n63n-6-$\ell$.
Claim 4.7.
The surface $Y$ is not $(\ell+\epsilon-1)$–defective [^1] Recall that a (non–degenerate, irreducible, projective) linearly normal surface $S\subset\mathbb{P}^{r}$ is said to be kk–defective if its kk–secant variety has dimension smaller than expected, or, equivalently, if imposing $k+1$ general double points to $|\mathcal{O}_{S}(1)|$ imposes less than $3(k+1)$ independent conditions..
Proof of Claim 4.7.
The list of kk–defective surfaces can be found in [3, Thm. 1.3], where they are divided in two types (i) and (ii). It is clear that $Y$ does not belong to case (i) for numerical reasons, so we have to exclude that $Y$ falls in case (ii). A kk–defective, non–degenerate surface in $\mathbb{P}^{r}$ in case (ii) is contained in a $(s+2)$-dimensional cone $W$ over a curve, with vertex a linear space $\Pi$ of dimension $s\leqslant k-1$ with $r\geqslant 2k+s+3$. We have to exclude that $Y$ is of this type, with $k=\ell+\epsilon-1$.
The surface $Y$ can be seen as the image of $\mathbb{F}_{1}$ via a linear system that we will soon identify. Let us denote by $E$ the (1)(-1)–curve of $\mathbb{F}_{1}$ and by $F$ a curve in the ruling of $\mathbb{F}_{1}$ over $\mathbb{P}^{1}$. The effective cone of $\mathbb{F}_{1}$ is spanned by $E$ and $F$. The linear system that maps $\mathbb{F}_{1}$ (isomorphically) to $Y$ is $|(3n-2-\ell)F+4E|$. One has
$((3n-2-\ell)F+4E)\cdot E=3n-6-\ell,\quad((3n-2-\ell)F+4E)\cdot F=4$
so that the minimum degree of a movable curve on $Y$ is $4$ (and it is achieved by the curves in $|F|$). One has
$\operatorname{deg}(Y)=(3n-2-\ell)^{2}-(3n-6-\ell)^{2}=4(3n-4+\epsilon).$
Suppose, by contradiction, that $Y$ is of type (ii) as above. Let $\mathfrak{R}$ be the (may be empty) 1–dimensional scheme cut out by the vertex $\Pi$ of the cone $W$ on $Y$, that has degree $\rho\geqslant 0$. Take a general hyperplane passing through $\Pi$. Such a hyperplane intersects $W$ in $\Pi$ plus $\gamma:=\operatorname{deg}(W)$ linear spaces of dimension $s+1$ containing $\Pi$, each of which intersects $Y$, off $\mathfrak{R}$, in $a\geqslant 1$ curves of degree $\delta\geqslant 4$. Looking at the section of such a hyperplane with $Y$, we see that
$4(3n-4+\epsilon)=\operatorname{deg}(X)=a\gamma\delta+\rho\geqslant 4\gamma,\quad\text{hence}\quad 3n-4+\epsilon\geqslant\gamma.$
On the other hand, one has $\gamma\geqslant r-s-1$, so that we find
$3n-4+\epsilon\geqslant r-s-1\geqslant r-k=15n-6\ell-\epsilon-15\geqslant 6n+2\epsilon-3$
that is clearly a contradiction. ∎
Claim 4.7 implies that
$\dim(\mathcal{P})=\frac{(3n-2-\ell)(3n+1-\ell)}{2}-\frac{(3n-6-\ell)(3n-5-\ell)}{2}-3\ell-3\epsilon=3n+\epsilon.$
Hence we see that in (10) equalities hold everywhere, and this proves that the $2\ell$ nodes of $\Sigma$ impose independent conditions to the linear system of surfaces of degree $n$ that contain the line $r$ with multiplicity n2n-2.
5. Deformations of surfaces with a double line
In this section we will consider Problem 1.1 for a surface $S$ of degree $n\geqslant 3$ that has only ordinary singularities along a line $R$. This means that: (a) the points of $R$ are double points for $S$, (b) at the general point $p\in R$ the tangent cone to $S$ at $p$ consists of two distinct planes containing $R$, (c) there are finitely many pinch points $p\in R$ such that the tangent cone to $S$ at $p$ consists of a double plane. A minimal resolution $f:S^{\prime}\longrightarrow S$ of $S$ is obtained by blowing–up $R$ in $\mathbb{P}^{3}$ and by taking the strict transform $S^{\prime}$ of $S$.
Again we will use a deformation argument that we are going to describe soon. The situation is similar to the one in Section 3.1, but slightly more complicated.
5.1. The ambient of the deformation
Let $\mathcal{X}^{\prime}=\mathbb{P}^{3}\times\mathbb{D}\longrightarrow\mathbb{D}$ be a trivial family parametrized by the disc $\mathbb{D}=\{t\in\mathbb{C}|\,|t|<1\}$, with fibres $\mathcal{X}^{\prime}_{t}\simeq\mathbb{P}^{3}$ over $t\in\mathbb{D}$.
Next we blow-up the line $R\subset\mathcal{X}^{\prime}_{0}$. Denote by $\mathcal{X}\longrightarrow\mathcal{X}^{\prime}$ the blowing-up of $\mathcal{X}^{\prime}$ along $R\subset\mathcal{X}^{\prime}_{0}$. Then $\mathcal{X}$ has fibre $\mathcal{X}_{t}\simeq\mathbb{P}^{3}$ if $t\neq 0$ and central fibre
$\mathcal{X}_{0}=A\cup\Theta,$
where $A$ is the blowing-up $p:A\longrightarrow\mathbb{P}^{3}$ of $\mathbb{P}^{3}$ along $R$ and $\Theta$ is the exceptional divisor, intersecting transversally $A$ along the smooth ruled surface
$E=A\cap\Theta\simeq\mathbb{P}(\mathcal{O}_{R}(1)\oplus\mathcal{O}_{R}(1))\cong\mathbb{F}_{0},$
coinciding on $A$ with the exceptional divisor of $p:A\longrightarrow\mathbb{P}^{3}$. We have
$\Theta\cong\mathbb{P}(\mathcal{O}_{R}(1)\oplus\mathcal{O}_{R}(1)\oplus\mathcal{O}_{R})\cong\mathbb{P}(\mathcal{O}_{R}\oplus\mathcal{O}_{R}\oplus\mathcal{O}_{R}(-1))$
that is a $\mathbb{P}^{2}$-bundle on $R$ and we denote by $P_{\Theta}\cong\mathbb{P}^{2}$ the linear equivalence class of a fibre of the natural morphism $\Theta\longrightarrow R$.
Lemma 5.1.
There is a natural isomorphism
$\Theta\cong{\rm Bl}_{R^{*}}({\mathbb{P}^{3}}^{*}),$
i.e., $\Theta$ can be seen in a natural way as the blowing-up of $\mathbb{P}^{3}$$^{*}$ along the dual line $R^{*}$ of $R$, namely $R^{*}$ is the pencil of planes containing $R$.
Proof.
This follows from [6, Prop. 9.11]. However we need to make the isomorphism $\Theta\cong{\rm Bl}_{R^{*}}({\mathbb{P}^{3}}^{*})$ very explicit.
Claim 5.2.
$E$ does not move on $\Theta$.
Proof of Claim 5.2.
Consider the obvious relation
$(A+\Theta)\cdot A\cdot\Theta=A^{2}\cdot\Theta+A\cdot\Theta^{2}=0,$
where $A^{2}\cdot\Theta=c_{1}(\mathcal{N}_{E|\Theta})$ and $A\cdot\Theta^{2}=c_{1}(\mathcal{N}_{E|A}):={\bf e}$. Denote by ${\bf f}_{E}$ and ${\bf r}_{E}$ the linear equivalence classes, respectively, of a fibre and a section of $E\cong\mathbb{P}^{1}\times\mathbb{P}^{1}$ as a $\mathbb{P}^{1}$-bundle over $R\simeq\mathbb{P}^{1}$, generating $\rm{Pic}(E)$ and such that ${\bf r}_{E}^{2}={\bf f}_{E}^{2}=0$.
By using that $K_{A}\sim p^{*}(K_{\mathbb{P}^{3}})+E$ and by the adjunction formula $\omega_{E}\simeq\omega_{A}\otimes\mathcal{N}_{E|A}$ one obtains that
$\mathcal{N}_{E|A}\sim\omega_{E}\otimes\omega_{{A}|_{E}}^{-1}\sim\omega_{E}\otimes(p^{*}(\mathcal{O}_{\mathbb{P}^{3}}(4))\otimes\mathcal{O}_{A}(-E))|_{E}.$
Since $\mathcal{O}_{A}(E)|_{E}={\bf e}$, we have
$2{\bf e}\sim-2{\bf f}_{E}-2{\bf r}_{E}+4{\bf f}_{E}=2{\bf f}_{E}-2{\bf r}_{E},$
hence
$c_{1}(\mathcal{N}_{E|A})={\bf e}\sim{\bf f}_{E}-{\bf r}_{E}\sim-\,c_{1}(\mathcal{N}_{E|\Theta}).$
This proves that $\mathcal{N}_{E|\Theta}$ is non-effective, as claimed.∎
Claim 5.3.
The linear system $|\mathcal{O}_{\Theta}(P_{\Theta}+E)|$ has dimension $3$ and its restriction to a plane $P\in|P_{\Theta}|$ coincides with the complete linear system $|\mathcal{O}_{P}(1)|$. Moreover, one has that $P_{\Theta}\cdot(P_{\Theta}+E)^{2}=1$, $(P_{\Theta}+E)|_{E}\sim{\bf r}_{E}$ and $E\cdot(P_{\Theta}+E)^{2}=0$.
Proof of Claim 5.3.
We first prove that $\mathcal{O}_{P}(P_{\Theta}+E)\simeq\mathcal{O}_{P}(E)\simeq\mathcal{O}_{P}(1).$ By the fact that $\Theta$ is rational, one has $h^{1}(\mathcal{O}_{\Theta})=h^{2}(\mathcal{O}_{\Theta})=0$. By the exact sequence
$0\longrightarrow\mathcal{O}_{\Theta}\longrightarrow\mathcal{O}_{\Theta}(P_{\Theta})\longrightarrow\mathcal{O}_{P}(P_{\Theta})\cong\mathcal{O}_{P}\longrightarrow 0,$
one sees that $h^{1}(\mathcal{O}_{\Theta}(P_{\Theta}))=0$. Twisting by $\mathcal{O}_{\Theta}(-P_{\Theta})$, one also obtains $h^{i}(\mathcal{O}_{\Theta}(-P_{\Theta}))=0,$ for $0\leqslant i\leqslant 3.$
Now, a fibre $P\sim P_{\Theta}$ of $\Theta$ cuts a fibre ${\bf f}_{E}\cong\mathbb{P}^{1}$ on $E$. Thus $P\cap E$ is a smooth rational plane curve. In order to see that $\mathcal{O}_{P}(E)\simeq\mathcal{O}_{P}(1)$, we prove that $h^{0}(\mathcal{O}_{P}(E))=3$. Consider the exact sequences
(11) $0\longrightarrow\mathcal{O}_{\Theta}(E-P_{\Theta})\longrightarrow\mathcal{O}_{\Theta}(E)\longrightarrow\mathcal{O}_{P}(E)\longrightarrow 0,$
(12) $0\longrightarrow\mathcal{O}_{\Theta}\longrightarrow\mathcal{O}_{\Theta}(E)\longrightarrow\mathcal{O}_{E}(E)\cong\mathcal{O}_{E}({\bf r}_{E}-{\bf f}_{E})\longrightarrow 0,$
and
(13) $0\longrightarrow\mathcal{O}_{E}({\bf r}_{E}-{\bf f}_{E})\longrightarrow\mathcal{O}_{E}({\bf r}_{E})\longrightarrow\mathcal{O}_{{\bf f}_{E}}({\bf r}_{E})\cong\mathcal{O}_{\mathbb{P}^{1}}(1)\longrightarrow 0.$
Using (13), one obtains $h^{1}(\mathcal{O}_{E}({\bf r}_{E}-{\bf f}_{E}))=0$. By (12) and by $h^{1}(\mathcal{O}_{\Theta})=0$, one has $h^{1}(\mathcal{O}_{\Theta}(E))=0$. It follows by (11) that
$h^{0}(\mathcal{O}_{P}(E))=h^{0}(\mathcal{O}_{\Theta}(E))+h^{1}(\mathcal{O}_{\Theta}(E-P_{\Theta})).$
By the exact sequence
$0\longrightarrow\mathcal{O}_{\Theta}(-P_{\Theta})\longrightarrow\mathcal{O}_{\Theta}(E-P_{\Theta})\longrightarrow\mathcal{O}_{E}(E-P_{\Theta})\cong\mathcal{O}_{E}({\bf r}_{E}-2{\bf f}_{E})\longrightarrow 0,$
using that $h^{i}(\mathcal{O}_{\Theta}(-P_{\Theta}))=0$, for $0\leqslant i\leqslant 3,$ one obtains that
$h^{1}(\mathcal{O}_{\Theta}(E-P_{\Theta}))=h^{1}(\mathcal{O}_{E}(E-P_{\Theta}))=h^{1}(\mathcal{O}_{E}({\bf r}_{E}-2{\bf f}_{E})).$
By Serre duality, $h^{1}(\mathcal{O}_{E}({\bf r}_{E}-2{\bf f}_{E}))=h^{1}(\mathcal{O}_{E}(-3{\bf r}_{E}))=2.$ Thus $h^{0}(\mathcal{O}_{P}(E))=3$, hence $\mathcal{O}_{P}(E)\simeq\mathcal{O}_{P}(1)$. Twisting (11) by $\mathcal{O}_{\Theta}(P_{\Theta})$, one obtains our first statement. The equality $P_{\Theta}\cdot(P_{\Theta}+E)^{2}=1$ follows. The relation $(P_{\Theta}+E)|_{E}\sim{\bf r}_{E}$ is trivial, hence $E\cdot(P_{\Theta}+E)^{2}=0$, as claimed. ∎
Now we can finish the proof of the lemma. First notice that the linear system $|P_{\Theta}+E|$ has no base points. Indeed, since it contains the linear system $|P_{\Theta}|+E$, the base points of $|P_{\Theta}+E|$ may occur only on $E$. But, by the exact sequence
$0\longrightarrow\mathcal{O}_{\Theta}(P_{\Theta})\longrightarrow\mathcal{O}_{\Theta}(P_{\Theta}+E)\longrightarrow\mathcal{O}_{E}(P_{\Theta}+E)\cong\mathcal{O}_{E}({\bf r}_{E})\longrightarrow 0,$
since, as we saw, $h^{1}(\mathcal{O}_{\Theta}(P_{\Theta}))=0$, $|P_{\Theta}+E|$ cuts out on $E$ the complete linear system $|{\bf r}_{E}|$ that is base point free.
From Claim 5.3 it follows that
$\phi_{|P_{\Theta}+E|}:\Theta\longrightarrow{\mathbb{P}^{3}}^{*}$
is a morphism that maps $\Theta$ birationally onto $\mathbb{P}^{3}$$^{*}$. The planes in $|P_{\Theta}|$ are mapped to planes in a pencil, that has as base locus a line $r$ where $E$ is contracted. The line $r$ can be naturally identified with $R^{*}$. Indeed, $\phi_{|P_{\Theta}+E|}$ contracts $E\cong\mathbb{F}_{0}$ along the ruling $|{\bf r}_{E}|$. Hence the image of $E$ can be identified with a fibre ${\bf f}_{E}$, that in turn can be identified with $R^{*}$. ∎
5.2. Deformation of a surface with a double line
Let now $S\subset\mathcal{X}^{\prime}_{0}$ be a degree $n\geqslant 3$ surface in $\mathbb{P}^{3}$ with only ordinary singularities along a line $R$. We denote by $H$ the Cartier divisor on $\mathcal{X}$ which is the pull-back of a general plane of $\mathbb{P}^{3}$ with respect to the projection morphism $\mathcal{X}\longrightarrow\mathbb{P}^{3}$. The strict transform $S_{A}$ of $S\subset\mathcal{X}^{\prime}_{0}$ on $\mathcal{X}$ sits on the component $A$ of $\mathcal{X}_{0}$, it coincides with the minimal desingularization of $S$ and it belongs to the linear system $|\mathcal{O}_{A}(nH-2E)|$. In particular, since $S$ has only ordinary singularities along $R$, then $S_{A}$ cuts on $E$ a smooth curve $C$.
Lemma 5.4.
The smooth curve $C$ on $E$ is linearly equivalent to
$C\sim n{\bf f}_{E}+2c_{1}(\mathcal{N}_{E|\Theta})\sim 2{\bf r}_{E}+(n-2){\bf f}_{E}$
hence it has genus n3n-3. The pencil $|{\bf f}_{E}|$ cuts out a $g^{1}_{2}$ on $C$ whose 2n42n-4 ramification points map to the pinch points of $S$ on $R$.
Proof.
We consider the line bundle $\mathcal{O}_{\mathcal{X}}(nH-2\Theta)$ whose restriction to the general fibre $\mathcal{X}_{t}\cong\mathbb{P}^{3}$, for $t\neq 0$, is $\mathcal{O}_{\mathbb{P}^{3}}(n)$. It restricts to $A$ as $\mathcal{O}_{A}(nH-2E)$. One has
$(nH-2\Theta)|_{\Theta}\sim(nH+2A)|_{\Theta}\sim nP_{\Theta}+2E$
and therefore $S_{A}\cdot E\sim n{\bf f}_{E}+2c_{1}(\mathcal{N}_{E|\Theta})\sim 2{\bf r}_{E}+(n-2){\bf f}_{E}$. The rest of the assertion is obvious. ∎
Let now $S_{0}:=S_{A}\cup S_{\Theta}\in|\mathcal{O}_{\mathcal{X}_{0}}(nH-2\Theta)|$, be a Cartier divisor with $S_{\Theta}\subset\Theta$. We want to understand what the surface $S_{\Theta}\subset\Theta$ can be, recalling that $\Theta$ can be seen as the blow–up of $\mathbb{P}^{3}$$^{*}$ along $R^{*}$ via the morphism determined by the linear system $|P_{\Theta}+E|$.
Set $H_{\Theta}\sim P_{\Theta}+E$. Since
(14) $(nH-2\Theta)|_{\Theta}\sim nP_{\Theta}+2E\sim nH_{\Theta}-(n-2)E$
the linear system $|(nH-2\Theta)|_{\Theta}|$ is sent by $\phi_{|P_{\Theta}+E|}$ to the linear system $\mathfrak{L}$ in $\mathbb{P}^{3}$$^{*}$ of surfaces of degree $n$ containing $R^{*}$ with multiplicity n2n-2. So $S_{\Theta}$ can be regarded as the normalization of a surface of degree $n$ in $\mathbb{P}^{3}$$^{*}$ with points of multiplicity n2n-2 along $R^{*}$, with the constraint that $S_{\Theta}$ cuts on $E$ the same curve $C$ that $S_{A}$ cuts on $E$, which, as we saw in Lemma 5.4, is a smooth curve in the linear system $|2{\bf r}_{E}+(n-2){\bf f}_{E}|$.
In what follows we will be interested in deformations of a surface of the type of $S_{0}$ in $\mathcal{X}$. If $S_{\Theta}$ is smooth, such a deformation can be interpreted as a total smoothing of $S$ in $\mathbb{P}^{3}$ (see [4, Prop. 3.2]).
5.3. The family of surfaces with a double line
In this section we only prove the following:
Proposition 5.5.
The family $\mathcal{F}_{n}$ of surfaces of degree $n$ in $\mathbb{P}^{3}$ that are singular along a line is irreducible, of dimension
$f_{n}={{n+2}\choose 3}+\frac{n(n-3)}{2}+3.$
The general surface in $\mathcal{F}_{n}$ has ordinary singularities along a line and no other singularity.
Proof.
Let $r\subset\mathbb{P}^{3}$ be a line, and let $\mathcal{S}_{n}$ be the linear system of surfaces of degree $n$ in $\mathbb{P}^{3}$ that are singular along $r$. It is easy to see, and can be left to the reader, that this linear system has dimension
$s_{n}={{n+2}\choose 3}+\frac{n(n-3)}{2}-1.$
Then the first assertion easily follows and the final assertion is clear. ∎
5.4. Nodal deformations of surfaces with a double line
In view of Problem 1.1, we are not interested in a total smoothing of $S$, but in deformations which are nodal, with possibly the maximum possible number of nodes. To obtain such a deformation, we will need to impose to $S_{\Theta}$ to have nodes off $E$, and possibly the maximum possible number of nodes. In order to be sure that we can preserve these nodes in a deformation off the central fiber, we should be able to apply [4, Thm. 4.6], so it will be necessary to know that the nodes of $S_{\Theta}$ impose independent conditions to the linear system of surfaces $|(nH-2\Theta)|_{\mathcal{X}_{0}}|$.
Recall that $\Theta$ can be seen as the blow–up of $\phi:=\phi_{|P_{\Theta}+E|}:\Theta\longrightarrow{\mathbb{P}^{3}}^{*}$ along $r:=R^{*}$ via the morphism determined by the linear system $|P_{\Theta}+E|$. As we saw in Section 5.2, the image $\Sigma$ of $S_{\Theta}$ via the map $\phi$ is a surface of degree $n$ in $\mathbb{P}^{3}$$^{*}$ that contains the line $r$ with multiplicity n2n-2. Suppose that $S_{\Theta}$ contains $\delta$ nodes. Then $\Sigma$ contains $\delta$ nodes off the line $r$.
Lemma 5.6.
Suppose that $S_{\Theta}$ contains $\delta$ nodes. Then they impose independent conditions to the linear system of surfaces $|(nH-2\Theta)|_{\mathcal{X}_{0}}|$ if and only if the $\delta$ nodes of $\Sigma$ impose independent conditions to the linear system of surfaces of degree $n$ in $\mathbb{P}^{3}$$^{*}$ passing through the line $r$ with multiplicity n2n-2.
Proof.
This is an immediate consequence of (14). ∎
Theorem 5.7.
Let $S\subset\mathbb{P}^{3}$ be a general surface of degree $n\geqslant 3$ that has only ordinary singularities along a line $R$. Then $S$ belongs to the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $\delta=3n-4-\epsilon$, where $0\leqslant\epsilon\leqslant 1$ has the same parity of $n$. In particular $\mathcal{F}_{n}$ is contained in $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$.
Proof.
Let us consider the situation described in Sections 5.1 and 5.2. In $\mathcal{X}_{0}$ we have the surface $S_{A}$ that we can identify as the minimal resolution $f:S^{\prime}\longrightarrow S$. Recall that $S_{A}$ cuts on $E=A\cap\Theta$ the smooth curve $C$. By Proposition 4.1, by the construction in Section 4.2, by Propositions 4.4 and 4.5, there is a surface $\Sigma$ of degree $n$ in $\mathbb{P}^{3}$ with mulptiplicity n2n-2 along a line $r$ and with $2\ell$ nodes (with $3n-4=2\ell+\epsilon$ and $0\leqslant\epsilon\leqslant 1$), such that, if $S^{\prime\prime}\longrightarrow\Sigma$ is the partial minimal desingularization of $\Sigma$ along $r$, then the pull–back of $r$ on $S^{\prime\prime}$ is the curve $C$. Then we can glue $S_{A}$ with $S_{\Theta}=S^{\prime\prime}$ along $C$, thus obtaining a Cartier divisor $S_{0}=S_{A}\cup S_{\Theta}\subset\mathcal{X}_{0}$. Now, taking into account Lemma 5.6 and the fact that the $2\ell$ nodes of $\Sigma$ impose independent conditions to the linear system of surfaces of degree $n$ that contain $r$ with multiplicity n2n-2 (see Proposition 4.5), we can apply [4, Thm. 4.6] and conclude the proof of the theorem.∎
The following remarks are in order.
Remark 5.8.
The proof of Theorem 5.7 shows that $S$, as in the statement of the theorem, can be deformed to a $\delta$–nodal surface (with $\delta=3n-4-\epsilon=2\ell$, where $0\leqslant\epsilon\leqslant 1$) in such a way that $\ell$ (arbitrary chosen) pinch points of $S$ deform each to a pair of nodes. This follows by the particular configurations of the nodes of the surface $S_{\Theta}$ (cf. Proposition 4.5 and Lemma 5.1).
Remark 5.9.
The codimension of the family $\mathcal{F}_{n}$ in $|\mathcal{O}_{\mathbb{P}^{3}}(n)|$ is
${{n+3}\choose 3}-1-f_{n}=3n-3.$
We proved that there is a regular component $V$ of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ contaning $\mathcal{F}_{n}$, and therefore $V$ has codimension $\delta=3n-4-\epsilon$. So, if $n$ is even, so that $\epsilon=0$, we see that $V$ has dimension just one more than $\mathcal{F}_{n}$. It follows that 3n43n-4 is the maximum $\delta$ such that $\mathcal{F}_{n}$ is contained in a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$. In the case $n$ odd we have $\delta=3n-5$, thus $V$ has dimension 2 more than $\mathcal{F}_{n}$ so that there is the possibility that $\mathcal{F}_{n}$ is contained in a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{3n-4}$. This is an interesting question to be investigated.